# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Interactive Environmental Learning in Physical Embodied Systems

### v4

@@ Abstract
Embodied agents face persistent challenges in real-world environments, including partial observability, limited spatial reasoning, and high-latency multi-memory integration. We present RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory under a parallelized architecture for efficient long-horizon planning and interactive environmental learning. A dynamic spatial knowledge graph (KG) ensures scalable and consistent memory updates, while a closed-loop planner with a critic module supports adaptive decision-making in dynamic settings. Experiments on EmbodiedBench show that RoboMemory, built on Qwen2.5-VL-72B-Ins, improves average success rates by 25% over its baseline and exceeds the closed-source state-of-the-art (SOTA) Gemini-1.5-Pro by 3%. Real-world trials further confirm its capacity for cumulative learning, with performance improving across repeated tasks. These results highlight RoboMemory as a scalable foundation for memory-augmented embodied intelligence, bridging the gap between cognitive neuroscience and robotic autonomy.
@@ 1 Introduction
Recent advances in Vision-Language Models (VLMs) (hurst2024gpt, ; bai2025qwen2, ) have enabled their growing use in embodied tasks (park2023generative, ; hu2023toward, ). VLM-based embodied agents can process multimodal inputs and generate high-level textual commands (e.g., “Pick up the cup”), which require translation via tool APIs to become executable robot actions. In contrast, Vision-Language-Action models (VLAs) (kim2024openvla, ; black2024pi_0, ; bjorck2025gr00t, ; chi2023diffusion, ) produce low-level control signals directly but generally rely only on the latest observation. This limits their ability to perform long-horizon, multi-step tasks that require reasoning over task history. In summary, VLA models enable direct robot control but lack high-level planning capabilities, and VLM-based embodied agents support strategic planning but struggle with direct motor control. This highlights a key gap inherent in two distinct technical approaches to embodied intelligence.
@@
To bridge this gap, recent work (yuan2025being, ; shi2025hi, ; tan2025roboos, ) proposes a “VLM planner + VLA executor” paradigm. Here, VLM-based embodied agents serve as high-level planners that decompose complex tasks (e.g., “make a coffee”) into executable sub-instructions (e.g., “grasp the cup”) that VLAs can complete. Although this paradigm improves performance on multi-step tasks, prior work suffers from two key limitations in real-world settings. First, real-world tasks (e.g., kitchen operations) require navigating across multiple locations to gather objects and tools, but the environment remains only partially observable at any time due to robots’ limited field of view and dynamic occlusions. This necessitates a planner with robust spatial awareness and long-term memory to maintain a consistent spatial awareness across viewpoints. However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs (yao2022react, )), which lack mechanisms for maintaining an overview of the environment’s spatial layout. Consequently, agents cannot reliably track object locations or recognize previously visited states. Second, pretrained VLMs are rarely trained on embodied planning trajectories, especially long-horizon, spatially grounded ones. So VLM-based agents often struggle to generalize to real-world settings (yang2025embodiedbench, ). To overcome these challenges, VLM-based planners must support interactive environmental learning — the ability to acquire, integrate, and retrieve spatial, episodic, and semantic knowledge during task execution, thereby enabling adaptation through experience.
To support long-horizon planning and interactive learning in real-world settings, agents require a comprehensive memory system with multiple specialized modules. Recent frameworks (tan2024cradle, ; glocker2025llm, ; wang2023voyager, ; agashe2024agent, ; fu2024msi, ; zhao2024expel, ; chen2024automanual, ) have integrated Retrieval-Augmented Generation (RAG)-based memory to enhance planning and interactive environmental learning (gao2023retrieval, ), but most are designed for simulated environments. A key limitation for real-world deployment is the absence of spatial memory, which is critical for building spatial awareness and providing context for planning. Additionally, existing multi-module memory systems often incur significant inference latency.
To overcome these limitations, especially the need for memory that is efficient, spatially grounded, and persistent in dynamic environments, we return to the essence of intelligence — how does the human brain plan, remember, and learn in dynamic environments? Inspired by cognitive neuroscience, we have designed RoboMemory, a parallel multi-memory architecture that simulates key functional regions of the brain. RoboMemory features a hierarchical and parallelized architecture enabling long-term planning and continuous adaptation. Drawing inspiration from cognitive neuroscience (milner1998cognitive, ), RoboMemory comprises four core components (Figure 1): (1) Information Preprocessor (thalamus-inspired) for multimodal sensory integration. (2) Comprehensive Embodied Memory System (hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (long-term, short-term, and sensory memory). Within this tiered system, four memory modules: Spatial, Temporal, Episodic, and Semantic operate under a unified, parallel-update paradigm to enable coherent knowledge integration while minimizing latency. (3) Closed-Loop Planning Module (prefrontal cortex-inspired) for high-level action sequencing. These three modules provide a high-level planner with comprehensive sensory and memorization ability. (4) Low-level Executor (cerebellum-inspired), consisting of a VLA-based operation model and a SLAM-based navigation model. The Low-level Executor directly controls the robot with low-level control signals to navigate and operate in the real-world environment.
To verify whether RoboMemory truly addresses the problems of long-horizon planning and interactive learning, we evaluate RoboMemory on EmbodiedBench, a long-horizon planning benchmark (yang2025embodiedbench, ). Using Qwen2.5-VL-72B as the base model, RoboMemory improves average success rates by 25% over its base model and 3% over the closed-source state-of-the-art model, Gemini-1.5-Pro. In real-world trials, RoboMemory executed diverse tasks twice consecutively: once for environmental familiarization (learning phase) and once for memory-augmented execution (testing phase) without resetting memory. The observed performance improvement validates RoboMemory’s capacity for interactive environmental learning. We further conduct ablation studies and error analysis to quantify component contributions and identify remaining limitations. We summarize our contribution as follows:
@@
  • We propose a brain-inspired unified embodied memory system, integrating four concurrently updated modules (Spatial, Temporal, Episodic, Semantic) into a single framework. It enables efficient, comprehensive memory operations and coherent knowledge integration, which are critical for interactive environmental learning in real-world embodied scenarios.
@@
  • We design a retrieval-based incremental update algorithm for real-time evolution of Spatial Knowledge Graphs (KGs). By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, it ensures efficient, consistent KG maintenance and addresses the scalability bottleneck of previous KG-based methods in embodied settings.
@@
  • RoboMemory supports interactive environmental learning for real-world physical robots: it enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
2 RoboMemory
RoboMemory is a hierarchical embodied agent system that equips robots with three core memory capabilities: historical interaction logs, dynamically updated spatial layouts, and accumulated task knowledge. As illustrated in Figure 2, each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
First, the information preprocessor converts multimodal sensor inputs into a textual summary of the current scene, which serves as the primary input to the Comprehensive Embodied Memory. Next, the Comprehensive Embodied Memory updates its internal representations, including action histories, object locations, and experiential knowledge. After information is updated, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module. Then, leveraging this contextual memory, the Closed-Loop Planning Module generates high-level, text-based action instructions. Finally, these commands are dispatched to low-level executors, who will directly control the robot and complete the instructions.
@@ 2 RoboMemory
Figure 2: RoboMemory architecture. (a) Left: Parallel Step Summarizer and Query Generator generate updates/queries for Comprehensive Embodied Memory. These memories enable Closed-Loop Planning for tasks like “slice and pick up the apple”—the Planner generates plans, while the Critic and memories adjust decisions via feedback from visual inputs/results. (b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic. Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
2.1 Information Preprocessor
At each time step $i$, RoboMemory receives a visual observation $\mathcal{O}_{i}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations.
@@ 2.1 Information Preprocessor
Since raw visual data is unsuitable for direct use in memory construct and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules. The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: Transforms $\mathcal{O}_{i}$ into a concise textual description $s_{i}$ of the just-executed action. The string $s_{i}$ is stored in the system’s working memory. (2) Query generator $\mathcal{Q}$: Derives a query $q_{i}$ from the same observation $\mathcal{O}_{i}$, which is used to probe long-term memory for relevant episodes.
@@
2.2 Comprehensive Embodied Memory System
We present a high-level memory system designed to enhance embodied agents’ long-horizon planning and enable interactive learning in dynamic, partially observable environments. Built on RAG and KG technologies (gutierrez2024hipporag, ), the system comprises Spatial-Temporal Memory and Long-Term Memory. Spatial-Temporal Memory continuously encodes object relationships and environmental layouts, as well as detailed task execution traces. These records inform Long-Term Memory, which goes through spatial-temporal information collected in each task, capturing what strategies succeeded or failed and relevant environmental facts (Figure 2).
The Comprehensive Embodied Memory module receives summarized sensory information from the information preprocessor to update both memory components in real time. Critically, all updates and retrievals are performed in parallel, ensuring efficient operation even under multiple active memory modules.
2.2.1 Spatial-Temporal Memory System
Due to the dynamic and partially observable nature of real embodied environments, completing long-term tasks requires dynamically maintaining a spatial-temporal memory system. The system records the agent’s interaction history with the environment and maintains a continuously updated spatial model of the current surroundings. To achieve these two purposes, we design spatial memory and temporal memory. The combination of spatial and temporal memory provides the embodied agent with comprehensive contextual awareness, supporting effective planning and interactive environmental learning in real-world settings.
@@ 2.2.1 Spatial-Temporal Memory System
Temporal Memory. Interactions between the robot and the environment must be recorded. For such temporally sequential memory, a simple structure suffices: a sequential buffer with automatic summarization triggered upon reaching capacity. The buffer tracks the high-level actions from the Embodied Agent, how the low-level executor actually controls the robot, and the environmental feedback received. In a specific design, temporal memory can store up to $N$ interaction summaries, each generated by an information preprocessor. When the buffer is full, rather than discarding old entries outright, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth.
@@
Spatial Memory. Current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds (zhang2023building, ; chang2023goat, ). However, such representations are overly detailed for high-level planning in embodied agents — precise geometric relationships (e.g., exact distances between objects) are unnecessary. Moreover, point clouds are difficult to update efficiently in dynamic scenes.
@@
To address these problems, we advocate for an abstracted spatial memory that captures object-level entities and their spatial relationships. Inspired by gutierrez2024hipporag , we represent this memory as a KG: objects become vertices, and spatial relations are encoded as edges. The KG focuses on high-level spatial relations (e.g., “cup on table”, “key left of drawer”). Unlike geometric representations that encode precise coordinates, which are often irrelevant for high-level planning, our KG focuses on semantically meaningful, task-relevant relations. This abstraction enhances the agent’s spatial reasoning capability in dynamic environments. In our memory system, spatial memory is a dynamic KG storing high-level spatial knowledge. To overcome the limited spatial reasoning of embodied agents in dynamic environments, we use a KG.
@@
We introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory. As illustrated in the right panel of Figure 2, the update process proceeds in four steps: (1) retrieves the most relevant sub-KG around new observations. (2) Injects new relations from the current observation by a VLM-based Relation Retriever. (3) Detects and resolves conflicts between newly extracted relations and existing ones (e.g., “cup on table” vs. ‘cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges. (4) Merges back and prunes isolated vertices. Moreover, our retrieval-based incremental update algorithm comes with provable efficiency guarantees: for a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix C.3 for formal analysis). Further architectural and implementation details are provided in Appendix C.1.
2.2.2 Interactive Environmental Learning System
Although episodic memory can record “what happened,” it cannot directly answer “why it succeeded/failed” or “what should be done next time.” Drawing on cognitive psychology’s classification of human long-term memory, we divide our system into episodic and semantic memory: the former records agent-environment interaction histories, while the latter extracts experiential insights to support long-term task reasoning. This update process mirrors human daily experience consolidation during sleep (maboudi2024retuning, ).
Episodic Memory: It captures task-level interactions, accounting for temporal interdependencies between sequential tasks in the same environment. The agent must memorize what it has done before to complete future tasks. Moreover, task-level interactions can be a reference that may help the agent improve its plan in the future.
Semantic Memory: It accumulates step-by-step action usage experiences (based on invoked actions and outcomes) to inform action arrangement. Post-task, it summarizes temporal memory, distilling successes from completed tasks and identifying failure causes/improvement strategies from unsuccessful ones, thus enabling both action-level and task-level learning across both task and action levels.
In implementation, both episodic and semantic memory share the same RAG framework consisting of an extractor, updater, and RAG storage (each entry is a memory entity). Post-task, the Extractor summarizes the task’s Spatial-Temporal Memory into a new memory entity. The RAG then retrieves similar existing entities (old information) from the RAG. Then, the Updater deletes, adds, or updates old memory entities according to new information. After that, we write the updated memory entities back to the RAG. Because we only update memory entities similar to new information, efficiency is ensured by restricting updates to old memory entities relevant to the new entity instead of all memory entities stored in the RAG.
2.3 Closed-Loop Planning Module for Dynamic Environment

### v6

@@ Abstract
Embodied intelligence aims to enable robots to learn, reason, and generalize robustly across complex real-world environments. However, existing approaches often struggle with partial observability, fragmented spatial reasoning, and inefficient integration of heterogeneous memories, limiting their capacity for long-horizon adaptation. To address this, we introduce RoboMemory, a brain-inspired framework that unifies Spatial, Temporal, Episodic, and Semantic memory within a parallelized architecture for efficient long-horizon planning and interactive learning. Its core innovations are a dynamic spatial knowledge graph for scalable, consistent memory updates and a closed-loop planner with a critic module for adaptive decision-making. Extensive experiments on EmbodiedBench show that RoboMemory, instantiated with Qwen2.5-VL-72B-Ins, improves the average success rate by 26.5% over its strong baseline and even surpasses the closed-source SOTA, Claude-3.5-Sonnet. Real-world trials further confirm its capability for cumulative learning, with performance consistently improving over repeated tasks. Our results position RoboMemory as a scalable foundation for memory-augmented embodied agents, bridging insights from cognitive neuroscience with practical robotic autonomy.
@@ I Introduction
Recent advances in Vision-Language Models (VLMs) [23, 6] have enabled their growing use in embodied tasks [33, 22]. VLM-based embodied agents can process multimodal inputs and generate high-level textual commands (e.g., “Pick up the cup”), which require translation via tool APIs to become executable robot actions. In contrast, Vision-Language-Action models (VLAs) [24, 8, 7, 13] produce low-level control signals directly but generally rely only on the latest observation. This limits their ability to perform long-horizon, multi-step tasks that require reasoning over task history. In summary, VLA models enable direct robot control but lack high-level planning capabilities, and VLM-based embodied agents support strategic planning but struggle with direct motor control. This highlights a key gap inherent in two distinct technical approaches to embodied intelligence.
@@
To bridge this gap, recent work [50, 34, 38] proposes a “VLM planner + VLA executor” paradigm. Here, VLM-based embodied agents serve as high-level planners that decompose complex tasks (e.g., “make a coffee”) into executable sub-instructions (e.g., “grasp the cup”) that VLAs can complete. Although this paradigm improves performance on multi-step tasks, prior work suffers from two key limitations in real-world settings. First, real-world tasks (e.g., kitchen operations) require navigating across multiple locations to gather objects and tools, but the environment remains only partially observable at any time due to robots’ limited field of view and dynamic occlusions. This necessitates a planner with robust spatial awareness and long-term memory to maintain a consistent spatial awareness across viewpoints. However, most VLM-based agents rely on chat-style context windows (e.g., logging instruction – feedback pairs [49]), which lack mechanisms for maintaining an overview of the environment’s spatial layout. Consequently, agents cannot reliably track object locations or recognize previously visited states. Second, pretrained VLMs are rarely trained on embodied planning trajectories, especially long-horizon, spatially grounded ones. So VLM-based agents often struggle to generalize to real-world settings [47]. To overcome these challenges, VLM-based planners must support interactive environmental learning — the ability to acquire, integrate, and retrieve spatial, episodic, and semantic knowledge during task execution, thereby enabling adaptation through experience. In summary, the current VLM–VLA paradigm still lacks (1) a robust spatial and long-term memory mechanism to maintain a consistent view of the environment, and (2) an effective way to acquire and leverage embodied planning experience for generalization.
To address these challenges, we propose RoboMemory, a brain-inspired memory framework designed to bridge the gap between biological cognitive mechanisms and robotic embodied intelligence. Drawing on foundational cognitive theories, we recognize that human memory is inherently organized into a tiered architecture rather than a flat buffer: sensory memory captures transient perceptual inputs [37], short-term working memory handles immediate information manipulation [4, 5], and long-term memory provides persistent storage for episodic events and semantic facts [3, 43]. Neuroscientific research further reveals that specialized brain regions underpin these functions: the thalamus integrates multimodal sensory inputs, the hippocampus consolidates spatial and episodic experiences into long-term storage [29], the prefrontal cortex regulates high-level planning based on retrieval, and the cerebellum coordinates low-level motor execution. Mimicking this biological hierarchy to ensure layered information processing, our recipe for this generalist memory system consists of four key stages (Figure 1): 1) Information Preprocessing (Thalamus-inspired), which integrates multimodal sensory inputs for downstream processing; 2) Comprehensive Embodied Memory (Hippocampus-inspired), which organizes experiential and spatial knowledge through a three-tier structure (sensory, short-term, and long-term). Crucially, we design four parallel-update modules (Spatial, Temporal, Episodic, and Semantic) to minimize latency while ensuring data consistency; 3) Closed-Loop Planning (Prefrontal Cortex-inspired), where the planner utilizes retrieved memory to generate high-level action sequences; 4) Low-level Execution (Cerebellum-inspired), which translates plans into precise robot actions using VLA models and SLAM-based navigation.
While recent frameworks [39, 18, 45, 1, 16, 54, 10] have integrated Retrieval-Augmented Generation (RAG) to enhance planning, they suffer from a critical limitation: most are tailored for simulated environments or rely on flat, text-based retrieval that fails to capture the spatial complexity of the physical world. These approaches often lack a hierarchical structure, leading to inefficiencies in long-term knowledge accumulation. Moreover, existing multi-module systems [38, 1] typically incur significant inference latency due to sequential memory updates. In contrast, our framework leverages a parallel-update paradigm and a hierarchical organization, enabling the agent to maintain a dynamic, spatially-aware worldview in real-time, which is essential for robust interaction in dynamic environments.
We evaluate RoboMemory on EmbodiedBench, a challenging long-horizon planning benchmark [47]. Using Qwen2.5-VL-72B as the base model, RoboMemory achieves state-of-the-art results, improving average success rates by 26.5% over the baseline and outperforming the closed-source Claude3.5-Sonnet [2]. In real-world deployments, our system demonstrates the capacity for continuous improvement: by executing diverse tasks sequentially without memory resets, the agent effectively learns from environmental familiarization.
Our main contributions are three-fold:
@@
  • We propose RoboMemory, a brain-inspired, unified embodied memory system that integrates four parallel update modules into a single framework. It enables efficient, comprehensive memory operations and coherent knowledge integration, which are critical for interactive environmental learning in real-world embodied scenarios.
@@
  • We design a retrieval-based incremental update algorithm for the real-time evolution of Spatial Knowledge Graphs (KGs). This algorithm bridges the gap between static knowledge representation and dynamic embodied interaction, establishing KGs as a practical and scalable modality for Spatial Memory. By retrieving relevant subgraphs, detecting local inconsistencies, and merging new observations, our method ensures efficient maintenance and overcomes the scalability bottlenecks that previously hindered KG-based approaches.
@@
  • RoboMemory supports interactive environmental learning in real-world environments. It enables sequential diverse tasks without memory reset, with experience accumulation driving steady performance improvements, demonstrating practical long-term autonomous learning in physical scenarios.
@@
Figure 2: (a) Left: The loop where the Planner, Critic, and Comprehensive Embodied Memory interact to adjust plans based on real-time visual inputs. Colored text denotes the execution status of actions (success/rejected/replanned). (b) Right: Spatial memory maintains a relevance/similarity-updated KG, and Semantic/Episodic memory manages a Vector DB with analogous logic. Besides, Temporal memory is implemented as a linear FIFO buffer that stores step-wise summaries generated by the Step Summarizer.
II Related Work
II-A VLM/LLM-based Agentic Frameworks in Embodied Tasks
II-B Spatial Memory
Spatial information is critical for embodied agents. Prior work has explored three main representations, each with inherent limitations that motivate our high-level abstractions: 2D Semantic Maps encode environments as pixel grids [51]. As the semantic maps are flattened, they discard 3D relationships (e.g., “above/beneath”) and precise geometry, limiting expressiveness for manipulation tasks requiring fine-grained spatial reasoning. 3D Point Clouds preserve accurate 3D information [52]. However, a 3D point cloud only encodes low-level spatial information. We need to put in extra effort to help agents understand 3D Point Clouds. Semantic Scene Graphs [38, 28, 19, 48, 20] offering a compact, high-level representation directly amenable to planning. However, existing methods suffer from two key drawbacks: (1) Most previous methods adopt rigid tree-like structures, which are easy to update with new information, but cannot model non-hierarchical or cyclic spatial configurations. (2) Some previous methods try to use general graphs to store various spatial information. [20] However, their design often lacks efficient dynamic update mechanisms—adding or modifying relations typically requires expensive full-graph recomputation, creating a scalability bottleneck for real-time embodied interaction.
For clarity, we summarize the differences among different memory systems in the previous work. The comparison is shown in Table III in the appendix. Additionally, we provide more ablation studies in Appendix -A.
III RoboMemory
RoboMemory is a hierarchical embodied agent system that equips robots with three core memory capabilities: historical interaction logs, dynamically updated spatial layouts, and accumulated task knowledge. As illustrated in Figure 2, in each iteration, RoboMemory follows a process of “Perception – Memory – Retrieval – Planning – Execution” process, ensuring that the agent continuously calibrates its memory and behavior in dynamic environments.
First, the information preprocessor converts multimodal sensor inputs into a textual summary of the current scene, which serves as the primary input to the Comprehensive Embodied Memory. Next, the Comprehensive Embodied Memory updates its internal representations, including action histories, object locations, and experiential knowledge. After updating the information, the memory system retrieves contextually relevant entries to inform the Closed-Loop Planning Module. Then, leveraging this contextual memory, the Closed-Loop Planning Module generates high-level, text-based action instructions. Finally, these commands are dispatched to low-level executors, which directly control the robot and complete the instructions. The execution process is demonstrated in Algorithm 1 in the appendix.
III-A Information Preprocessor
@@
At each time step $t$, RoboMemory receives a visual observation $\mathcal{O}_{t}$: an RGB frame (in simulation) or a short video clip (on physical robots), representing the agent’s observations. Since raw visual data is unsuitable for direct use in memory construction and retrieval, RoboMemory first employs an information preprocessor to convert multimodal observations into textual representations, thereby providing a semantic interface for subsequent memory and planning modules. The information preprocessor executes two Vision-Language Models (VLMs) in parallel: (1) Step summarizer $\mathcal{S}$: It transforms $\mathcal{O}_{t}$ into a concise textual description $s_{t}$ of the just-executed action. The string $s_{t}$ is stored in the system’s working memory. (2) Query generator $\mathcal{Q}$: It derives a list of queries $q_{t}=[q_{t}^{(1)},q_{t}^{(2)},\dots,q_{t}^{(N)}]$ from the same observation $\mathcal{O}_{t}$. Each query $q^{(i)}_{t}$ is a natural language-based query. These queries are used to query information from the memory system that may be useful.
@@
III-B Comprehensive Embodied Memory
To address the long-term memory limitations in current embodied agent frameworks, we propose the Comprehensive Embodied Memory System. This system consists of multiple memory modules. We denote the memory system containing $L$ distinct modules as $M_{t}=[M_{t}^{(1)},M_{t}^{(2)},...,M_{t}^{(L)}]$, where $M_{t}$ represents the memory stored at step $t$, and $M_{t}^{(l)}$ denotes the $l$-th memory module. Generally, the update and retrieval process at iteration $t$ of a general memory is shown below:
$M_{t}=\mathcal{U}(M_{t-1},s_{t})$ (1)
$r_{t}=\mathcal{R}(M_{t},q_{t})$ (2)
First, we update memory modules with memory update algorithm $\mathcal{U}$, where we update the previous memory $M_{t-1}$ using the latest summarization $s_{t}$. Then, with updated $M_{t}$, we use memory retrieve algorithm $\mathcal{R}$ to retrieve the information that is useful for the planning module. In $\mathcal{R}$, we use queries $q_{t}$ to query $M_{t}$, yielding retrieval results from each module: $r_{t}=[r_{t}^{(1)},r_{t}^{(2)},...,r_{t}^{(L)}]$. These results are then passed to the planning module, which helps it plan future movements. However, sequentially updating and retrieving from $L$ modules would be a slow process. Therefore, we parallelize these steps across all modules, which significantly enhance the efficiency of the memory system.
The memory system consists of four distinct modules ($L=4$): Temporal Memory, Spatial Memory, Semantic Memory, and Episodic Memory. For efficiency, all memory modules are updated and retrieved in parallel. Thus, even with multiple modules, the system remains highly efficient. Functionally, inspired by cognitive psychology [27], our modules handle memory at different levels. In cognitive psychology, memory is divided into Sensory Memory, Short-term Memory, and Long-term Memory. Mirroring this hierarchy, our modules are organized as follows. First of all, the step summarizer $\mathcal{S}$ summarizes the agent’s interactions with the environment at each iteration. It acts as Sensory Memory. Secondly, Temporal Memory and Spatial Memory function as short-term memory. These two memories will be updated at each iteration. They are designed to store the information of sensory memory in every iteration. For Temporal Memory, we record the agent’s action history sequentially, while for Spatial Memory, we dynamically record the spatial relationships between different objects in the environment based on Sensory Memory. These memories can provide a relatively long and detailed history of the current task for the Closed-Loop Planner to make a future plan. Thirdly, Semantic Memory and Episodic Memory serve as Long-term Memory. They are updated only when meaningful information arises (e.g., after task completion). These memories store highly abstract knowledge, not limited to the current task, but synthesized from past experiences. This knowledge—factual, event-based, and experiential—improves the agent’s future task performance. It is the source of RoboMemory’s interactive learning capability. We now detail each module.
@@ III-B Comprehensive Embodied Memory
Temporal Memory. In the Temporal Memory, we record interactions between the robot and the environment (i.e., Sensory Memory) of each iteration sequentially. This information can provide the embodied agent with simple awareness of “What I have done”. For such temporally sequential memory, a simple structure is sufficient: a sequential buffer with automatic summarization triggered when the record sequence reaches its capacity. In a specific design, temporal memory can store up to $N$ interaction summaries, each generated by an information preprocessor. When the buffer is full, we compress the oldest $N$ steps into a single summarized entry using a VLM, which is then reinserted at the front of the buffer, ensuring continuous context retention without unbounded growth. However, the information from previous memories will gradually be lost as we summarize it multiple times. For retrieval, we provide all existing memory in text to downstream modules.
@@
Spatial Memory. The spatial memory is designed to dynamically record the high-level spatial relationships of different entities in the environment. However, current spatial memory approaches often rely on RGB-D cameras to reconstruct 3D point clouds [51, 9]. These representations are too detailed for high-level planning in embodied agents. For example, precise geometric relationships (e.g., exact distances between objects) are unnecessary.
@@
To address these problems, we use a dynamic KG to store high-level spatial information: objects and positions in the environment become vertices of the KG, and spatial relations between objects or positions are encoded as edges. The KG focuses on high-level spatial relations (e.g., “cup on table”, “key left of drawer”). By these settings, spatial KG focuses on semantically meaningful, task-relevant relations. This spatial information enhances the agent’s spatial reasoning capability in dynamic environments.
@@
However, as related work shows, most KG construction algorithms are designed for static long content. This does not meet the demand of using KG as spatial memory for the agent. The KG needs to update efficiently in response to new information. To address this issue, we introduce a retrieval-driven, incremental KG update algorithm that maintains a locally modifiable, globally consistent, and dynamically adaptive spatial memory. As illustrated in the right panel of Figure 2, the update process proceeds in four steps: (1) retrieves the most relevant sub-KG around new observations. (2) Injects new relations from the current observation by a VLM-based Relation Retriever. (3) Detects and resolves conflicts between newly extracted relations and existing ones (e.g., “cup on table” vs. “cup in drawer”) using a VLM-based resolver, which decides whether to add, delete, or modify edges. (4) Merges back and prunes isolated vertices. Moreover, our retrieval-based incremental update algorithm is accompanied by provable efficiency guarantees. For a KG with $n$ vertices and maximum degree $D$, the number of vertices processed per update is bounded by $O(D^{K})$, where $K$ is the retrieval hop distance (see Appendix -D for analysis). Further implementation details are provided in Appendix -C1.
Semantic Memory. In cognitive psychology, semantic memory stores time-independent facts. These facts are stable, update slowly, and require long-term retention. In RoboMemory, semantic memory records task-relevant experiences and environmental knowledge during execution. This information can help RoboMemory adapt to new environments or tasks. This information is highly abstract and does not need to be updated frequently. In RoboMemory, semantic memory updates when new information is encountered during execution, for example, after completing a subtask or encountering important information. To store and update memory efficiently, we design a memory management system based on a vector database (DB). In the vector DB, each experience/fact is described in natural language (denoted as a memory item). Each memory item is converted into a semantic vector for querying. For dynamic updates, we adapt a basic method from prior work [12]. However, we adopt it into the embodied environment. As shown in the bottom-right of Figure 2, the semantic memory update algorithm involves two VLM-based modules. Firstly, the Long-term Memory Creator generates new memory items based on short-term memory. We retrieve the top-$S$ most similar existing memory items from existing memory via cosine similarity. A VLM-based updater then compares new and existing items to decide whether to: add the new item, update an existing item, remove an outdated item, or perform Noop (if redundant). Since updates only involve a maximum of $S$ previous memory items and the update process is parallelized across all memory modules, this update method ensures that semantic memory remains efficient even as the database grows. For retrieval, we use the same process as a traditional vector DB. We use queries to extract top-$N$ relevant information for downstream modules.
Episodic Memory. In cognitive psychology, Episodic Memory is another important part of long-term memory. It can store task-specific execution summaries (i.e., “autobiographical” records of past attempts). In RoboMemory, the Episodic Memory module is responsible for recording every interaction trajectory it has gone through, including the sequence of actions the robot did and the feedback from the environment. The trajectory information can help to improve the planning ability of RoboMemory. For example, if a trajectory for completing a similar task is stored in episodic memory, it can guide the agent in completing the current task. As the agent only needs to follow the successful trajectory in the memory, it can reduce hallucinations or errors in the VLM planner. Technically, Episodic Memory shares the same storage and VLM-driven vector DB update mechanism as Semantic Memory, ensuring consistent architectural design.
III-C Closed-Loop Planning Module